# Back-Reaction of Super-Hubble Fluctuations, Late Time Tracking and Recent Observational Results

### v1

@@ Abstract
Previous studies suggested that the back-reaction of super-Hubble cosmological fluctuations could lead to a dynamical relaxation of the cosmological constant. Moreover, this mechanism appears to be self-regulatory, potentially leading to an oscillatory behavior in the effective dark energy. Such an effect would occur in any cosmological model with super-Hubble matter fluctuations, including the standard $\Lambda$CDM model, without the need for additional mechanisms.
The recent DESI data, which supports the possibility of dynamical dark energy, has renewed interest in exploring scenarios leading to such oscillatory behavior. In this work we propose a parametrization to account for the impact of super-Hubble fluctuations on the background energy density of the Universe. We model the total effective cosmological constant as the sum of a constant piece and an oscillating contribution. We perform a preliminary comparison of the background dynamics of this model with the recent radial BAO data from DESI. We also discuss the status of the $H_{0}$ tension problem in this model. While this analysis serves as a first step in testing the scenario, we emphasize the importance of considering this effect, as it is naturally expected to be present and could offer valuable insights in light of the recent observational data.
@@ I Introduction
In the standard framework of cosmology, the $\Lambda$CDM model, the Dark Energy (DE) component, which comprises the majority of today’s cosmological energy density, is parametrized by a positive cosmological constant term, $\Lambda$, in the Einstein equations. Despite its apparent simplicity, this interpretation faces conceptual problems. In addition to some well known problems, such as the cosmological constant problem, the coincidence problem and the Hubble tension problem, several works in the literature have been pointing to an instability of a de Sitter phase. In particular, there are indications that a de Sitter space could be unstable due to infrared (IR) effects Polyakov:2012uc ; Polyakov:2007mm ; Mazur:1986et ; Mottola:1985qt ; Mottola:1984ar ; Tsamis:1994ca  [^1] In the context of String Theoriy there have also been indications that it is not possible to obtaine a stable de Sitter phase Obied:2018sgi ; Dvali:2014gua ; Danielsson:2018ztv ; Dasgupta:2018rtp , and the recently suggested “Trans-Planckian Censorship Criterion” Bedroya:2019snp ; Bedroya:2019tba ; Brandenberger:2021pzy implies that a positive cosmological constant is not consistent from an effective field theory point of view.. Already in the work of Ref. Tsamis:1992sx ; Tsamis:1996qq it has been suggested that the reason that our universe is not inflating today is that IR processes in pure quantum gravity (no matter) tend to screen the bare $\Lambda$. However, the authors of Tsamis:1996qq found that these effects become important only after perturbation theory has broken down. In 1997, in the work of Ref.Abramo:1997hu the back-reaction of super-Hubble modes of cosmological perturbations induced by matter were studied, finding that the effective energy density associated with these modes would counteract any pre-existing $\Lambda$ (see also Finelli:2001bn ; Finelli:2003bp ; Marozzi:2006ky ). Later, in 1999, in the work of Brandenberger:1999su it was speculated that this dynamical relaxation mechanism for $\Lambda$ could be self-regulating  [^2] Note also the related “Everpresent Lambda” scenario, which was proposed by drawing on ideas from causal set theory and unimodular gravity, and which leads to an effective fluctuating cosmological “constant”, see Refs.Ahmed:2002mj ; Das:2023hbw ; Das:2023rvg , for instance. . A crucial question is whether this back-reaction effect can be locally measured Unruh:1998ic ; Geshnizjani:2002wp ; Abramo:2001dc . Recent works have shown that super-Hubble fluctuation modes can in fact modify the parameters of a local Friedmann cosmology in the context of a more realistic scenario with more than one matter field, admitting entropy fluctuations (in this context see the works of Refs Geshnizjani:2003cn ; Marozzi:2012tp ; Finelli:2011cw ; Gasperini:2009wp ; Gasperini:2009mu ; Marozzi:2010qz ; Marozzi:2011zb ; Abramo:2001dd ; Losic:2005vg ; Losic:2006ht ; Brandenberger:2004ix ; Afshordi:2000nr ; Comeau:2023gxk ; Comeau:2023euf )  [^3] In the context of the back-reaction of sub-Hubble cosmological fluctuations, a recent analysis can be found, for instance, in Ref.Giani:2024nnv . In the work of Brandenberger:2018fdd the analysis was extended beyond perturbation theory, showing that the locally measured Hubble expansion rate indeed obtains a negative contribution from the back-reaction of super-Hubble fluctuations, whose amplitude grows in time. This supports the suggestion that (quasi) de Sitter space-times are unstable, and that this back-reaction mechanism might lead to a dynamical relaxation of the cosmological constant. As first speculated in Ref.Brandenberger:1999su , this is expected to lead to an effective dark energy density which oscillates in time around the dominant matter density.
@@
The quest for understanding the nature and dynamics of the accelerated expansion of the Universe became even more interesting recently with the improvement in the measurements of the expansion history, which are increasingly allowing us to constrain the expansion history of the Universe. More recently, the Dark Energy Spectroscopic Instrument (DESI) Year 1 results were released DESI:2024uvr ; DESI:2024lzq , officially marking the start of the Stage IV DE era. The cosmological results stem from measurements of baryon acoustic oscillations (BAO) in galaxy, quasar and Lyman-$\alpha$ forest tracers. The DESI results, when combined with a number of external probes, provide intriguing hints for a dynamical, time-evolving DE component DESI:2024mwx ; DESI:2024aqx ; DESI:2024kob .  [^4] See, however, Payeur:2024dnq for caveats. These hints may have enormous implications for the understanding of the nature of DE.
@@
In light of these results, recent works, among them Jiang:2024xnu and Escamilla:2024fzq , have pointed to signals, from late-time cosmological data, of oscillatory/non-monotonic features in the shape of the expansion rate at low redshifts. The possibility of an oscillating dark energy has already been addressed in earlier works, as for instance in RefsZhao:2017cud ; Tamayo:2019gqj ; Rubano:2003er ; Linder:2005dw ; Feng:2004ff ; Nojiri:2006ww ; Kurek:2007bu ; Jain:2007fa ; Saez-Gomez:2008mkj ; Kurek:2008qt ; Pace:2011kb ; Pan:2017zoh ; Panotopoulos:2018sso ; Rezaei:2019roe ; Yao:2022jrw ; Rezaei:2024vtg ; Tian:2019enx DESI:2024kob ; Adil:2023exv ; Mbonye:2022cnf ; Mbonye:2024mss ]. In Ref.Escamilla:2024fzq , in light of the recent cosmological data, twelve different parametrizations for an oscillatory dark energy equation of state were analyzed, among them the ones proposed in Refs. Linder:2005dw ; Feng:2004ff ; Pan:2017zoh ; Zhao:2005vj . This analysis demonstrated that all the oscillating DE parametrizations considered provided a better fit to the data when compared to the cosmological constant, although in the Bayesian evidence analysis they are penalized due to the additional free parameter. In a similar context, in the work of Ref.Jiang:2024xnu , the late-time expansion history was reconstructed in a way, according to the authors, to be “as model-independent and non-parametric as possible”. In their first reconstruction, using DESI+PantheonPlus data, two features in the unnormalized expansion rate, were identified, a bump at redshift $z\sim 0.5$, and a depression at $z\sim 0.9$ (both features were absent when replacing the DESI dataset with the older SDSS one). In spite of the higher number of unique redshifts, the effective volume covered by DESI is still smaller than the one of SDSS, so whether the abovementioned features are real or due to sample variance fluctuations is something expected to be clarified with the upcoming releases from DESI. On the other hand, the depression feature was noticeable in both the DESI+PantheonPlus and DESI+Union3 reconstructions, but not in the DESI+DESY5 one  [^5] In addition to the upcoming releases from DESI, the presence of this feature can also be checked by future SNeIa samples.. For more information on the reconstruction methods see Ref.Jiang:2024xnu . While the Chevallier-Polarski-Linder (CPL) fit, the $w_{0}-w_{a}$ parametrization usually considered when interpreting DESI data, is able to partially capture one of these features, it cannot fit both, since the model does not have sufficient complexity. The authors of Jiang:2024xnu conclude that fitting both features requires an oscillatory/non-monotonic behavior of the Hubble expansion rate.
@@
While a more conclusive result will need to wait for more precise future data, we believe it is important to already now further investigate theoretical mechanisms that could account for a non-monotonic dynamical behavior of DE. As discussed above, the back-reaction effect of super-Hubble fluctuations is one mechanism that could naturally lead to an oscillatory behavior for the dark energy density, and consequently for the Hubble parameter, without the need for extra ingredients beyond the standard $\Lambda$CDM model. This back-reaction effect is present in any cosmological model whenever entropic super-Hubble cosmological fluctuations are present, not relying on any extra ingredients. We believe that, in light of the new observational results, it is important not to neglect the possible presence of such a mechanism.
@@
In the present work we propose a parametrization to account for the back-reaction effect of super-Hubble fluctuations by means of an oscillating effective energy density component. This may provide a full non-perturbative description of the back-reaction effect. We then analyze such a model considering data from the radial component of the Baryon Acoustic Oscillations from the recent DESI data. We compute the Hubble parameter at low redshifts as predicted by this model, and compare the result to the recent DESI data in order to estimate the maximum amplitude of the oscillations allowed by the data. We also analyze the predictions of the model for the effective dark energy equation of state, for the deceleration parameter, and for $H_{0}$.
@@
While our study can be considered to be a first-step analysis limited to background quantities, aiming only to highlight the importance of not neglecting the back-reaction mechanism, the scenario has the advantage of avoiding the need to make extra suppositions about a new matter source of DE. Nevertheless, in order to obtain more conclusive results, it is necessary to further analyze the predictions of the model, in particular taking into account the induced cosmological perturbations. This will allow us to test the model using additional BAO data beyond only the radial component, as well as using further data from other cosmological sources. This analysis is left for a future work.
@@ II A Model for the Back-Reaction of Super-Hubble fluctuations
In the absence of anisotropic stress, the scalar metric fluctuation variables $\phi$ and $\psi$ in eq.2 are equal. To write the metric, we have chosen a particular coordinate system (longitudinal gauge)  [^6] See e.g. Mukhanov:1990me for a review of the theory of cosmological perturbations.. In Eq. 2 $\gamma_{ij}$ is the background metric of the constant time hypersurfaces. We assume vanishing spatial curvature and then $\gamma_{ij}=\delta_{ij}$. In the above equations, the background metric is given by the scale factor $a(t)$ and the background matter by $\varphi_{0}(t)$. We will only consider here the effect of scalar metric fluctuations.
@@
where the first term on the right hand side of the equation is the contribution of the background matter field to $T_{\mu\nu}$. Improved studies of the back-reaction of super-Hubble fluctuations focus on physically measurable quantities such as the spatial average of the local Hubble expansion rate Geshnizjani:2003cn ; Marozzi:2012tp ; Finelli:2011cw ; Gasperini:2009wp ; Gasperini:2009mu ; Marozzi:2010qz ; Marozzi:2011zb ; Abramo:2001dd ; Losic:2005vg ; Losic:2006ht ; Brandenberger:2004ix ; Afshordi:2000nr , and even beyond perturbation theory Brandenberger:2018fdd . Some of these results have pointed to a self adjusting mechanism for the cosmological constant, as first discussed in Brandenberger:1999su .
The origin of the self-adjusting mechanism for the relaxation of the cosmological constant is as follows Brandenberger:1999su .: in an accelerating phase, fluctuation modes exit the Hubble radius and freeze out. Cosmological perturbations induced by super-Hubble matter fluctuations contribute to the effective energy-momentum tensor like a negative cosmological constant (negligible kinetic and tension energies leading to an effective equation of state $w=-1$, and negative effective energy density - negative since matter induces a potential well and the negativity of the gravitational energy overwhelms the positivity of the matter energy for super-Hubble modes). This will induce a decrease in the effective cosmological constant by counteracting the effect of the pre-existing $\Lambda$. But once the energy density $\rho_{DE}(t)$ in the effective cosmological constant has fallen below that of matter, accelerated expansion will stop, modes will re-enter the Hubble radius. The decrease in $\rho_{DE}$ will stop, and soon (within a fraction of the Hubble expansion time), $\rho_{DE}$ will again dominate, leading to a new period of acceleration which, in turn, leads to a renewed buildup of the back-reaction effect. This cycle implies an oscillatory behavior in the effective energy density associated to the back-reaction mechanism.
@@
where $A$ is a dimensionless amplitude, $\omega$ is the frequency of the oscillation, $\phi$ is a phase and $\rho_{m}$ is the background density of matter. The term $\omega t$ in the above equation can be written in terms of the redshift $z$ as $\omega t\equiv nz$, with $n$ being a constant, which we expect to have an order of magnitude not much smaller or bigger than one since the typical time scale of oscillations is expected to be set by the Hubble time. This ansatz expresses the idea that, after the time $t_{eq}$ of equal matter and radiation, the back-reaction term should track the energy density $\rho_{m}$ of matter. A priori it is expected that the amplitude $A(t)$ should be time-dependent. However, in the present work, we will focus in the late time predictions, since we are going to analyze our model in light of data from $2.50\leq z<$ 2.5. For this small redshift interval we believe it is reasonable to consider $A(t)\approx cte$ as a good approximation. Therefore in the rest of the paper we are going to consider the following expression for the effective density from the back-reaction effect, $\rho_{BR}\ =\,A\;cos(\omega t+\phi)\;\rho_{m}(t)$.
This contribution of the energy density from the back-reaction effect is added to the total energy of the universe, in such a way that the first Friedmann equation can be written as,
@@
Concerning the phase $\phi$, we will chose it in a way that the back-reaction mechanism begins at a redshift $z=3400$ with $\rho_{BR}=0$, when it starts building up. This choice is reasonable since we are taking the back-reaction energy density as being proportional to the matter energy density, and before recombination it would rather track the radiation component.
@@
$\begin{split}w_{BR}&=\left(\frac{H^{\prime}H}{a}-\frac{3H^{2}}{2}+\frac{3H_{0}^{2}\;\Omega_{\Lambda 0}}{2}\right)\frac{1}{4\pi G}\;\frac{1}{\rho_{BR}},\\ &=\frac{n\left(z+1\right)}{3\tan\left(n(z-3400)\right)}\end{split}$ (13)

### v2

@@ Abstract
Previous studies suggested that the back-reaction of super-Hubble cosmological fluctuations could lead to a dynamical relaxation of the cosmological constant. Moreover, this mechanism appears to be self-regulatory, potentially leading to an oscillatory behavior in the effective dark energy. Such an effect would occur in any cosmological model with super-Hubble matter fluctuations, including the standard $\Lambda$CDM model.
The recent DESI data, which indicate that dark energy may be dynamical, has renewed interest in exploring scenarios leading to such oscillatory behavior. In this work we propose a parametrization to account for the impact of super-Hubble fluctuations on the background energy density of the Universe. We model the total effective cosmological constant as the sum of a constant piece and an oscillating contribution. We perform a preliminary comparison of the background dynamics of this model with the recent radial BAO data from DESI. We also discuss the status of the $H_{0}$ tension problem in this model.
@@ I Introduction
In the standard framework of cosmology, the $\Lambda$CDM model, the dark energy (DE) component, which comprises the majority of today’s cosmological energy density, is parametrized by a positive cosmological constant term, $\Lambda$, in the Einstein equations. Despite its apparent simplicity, this interpretation faces conceptual problems. In addition to some well known problems, such as the cosmological constant problem, the coincidence problem and the Hubble tension problem, several works in the literature have been pointing to an instability of a de Sitter phase. In particular, there are indications that a de Sitter space could be unstable due to infrared (IR) effects Polyakov:2012uc ; Polyakov:2007mm ; Mazur:1986et ; Mottola:1985qt ; Mottola:1984ar ; Tsamis:1994ca  [^1] In the context of String Theory there have also been indications that it is not possible to obtain a stable de Sitter phase Obied:2018sgi ; Dvali:2014gua ; Danielsson:2018ztv ; Dasgupta:2018rtp , and the recently suggested “Trans-Planckian Censorship Criterion” Bedroya:2019snp ; Bedroya:2019tba ; Brandenberger:2021pzy implies that a positive cosmological constant is not consistent from an effective field theory point of view.. Already in the work of Ref. Tsamis:1992sx ; Tsamis:1996qq it has been suggested that the reason that our universe is not inflating today is that IR processes in pure quantum gravity (no matter) tend to screen the bare $\Lambda$. However, the authors of Tsamis:1996qq found that these effects become important only after perturbation theory has broken down. In 1997, the work of Ref.Abramo:1997hu studied the back-reaction of super-Hubble modes of cosmological perturbations induced by matter, finding that the effective energy density associated with these modes would counteract any pre-existing $\Lambda$ (see also Finelli:2001bn ; Finelli:2003bp ; Marozzi:2006ky ). Later, in 1999, in the work of Brandenberger:1999su it was speculated that this dynamical relaxation mechanism for $\Lambda$ could be self-regulating  [^2] Note also the related “Everpresent Lambda” scenario, which was proposed by drawing on ideas from causal set theory and unimodular gravity, and which leads to an effective fluctuating cosmological “constant”, see Refs.Ahmed:2002mj ; Das:2023hbw ; Das:2023rvg , for instance.. A crucial question that remained at that time was whether this back-reaction effect could be locally measured Unruh:1998ic ; Geshnizjani:2002wp ; Abramo:2001dc . Recent works have shown that super-Hubble fluctuation modes can in fact modify the parameters of a local Friedmann cosmology, when measured by a clockfield Geshnizjani:2003cn ; Marozzi:2012tp ; Finelli:2011cw ; Gasperini:2009wp ; Gasperini:2009mu ; Marozzi:2010qz ; Marozzi:2011zb ; Abramo:2001dd ; Losic:2005vg ; Losic:2006ht ; Brandenberger:2004ix ; Afshordi:2000nr ; Comeau:2023gxk ; Comeau:2023euf  [^3] In the context of the back-reaction of sub-Hubble cosmological fluctuations, a recent analysis can be found, for instance, in Ref.Giani:2024nnv  [^4] Another interesting approaches to account for the back-reaction effect can be seen in the works of Refs.Kolb:2009rp ; Kolb:2011zz ; Marra:2011ct , for instance.. In the work of Brandenberger:2018fdd the analysis was extended beyond perturbation theory, showing that the locally measured Hubble expansion rate indeed obtains a negative contribution from the back-reaction of super-Hubble fluctuations, whose amplitude grows in time. This supports the suggestion that (quasi) de Sitter space-times are unstable, and that this back-reaction mechanism might lead to a dynamical relaxation of the cosmological constant. As first speculated in Ref.Brandenberger:1999su , this is expected to lead to an effective dark energy density which oscillates in time around the dominant matter density.
@@
The quest for understanding the nature and dynamics of the accelerated expansion of the Universe became even more interesting recently with the improvement in the measurements of the expansion history, which are increasingly allowing us to constrain the expansion history of the Universe. Recently, the Dark Energy Spectroscopic Instrument (DESI) Year 1 (DR1)DESI:2024aqx ; DESI:2024kob ; DESI:2024mwx and DESI DR2 DESI:2025zgx ; DESI:2025fii results were released, officially marking the start of the Stage IV DE era. The cosmological results stem from measurements of baryon acoustic oscillations (BAO) in galaxy, quasar and Lyman-$\alpha$ forest tracers. The DESI results, when combined with a number of external probes, provide intriguing hints for a dynamical, time-evolving DE component DESI:2024mwx ; DESI:2024aqx ; DESI:2024kob . These hints were present already in the DR1 release  [^5] See, however, Payeur:2024dnq for caveats.. Later, the enhanced statistical power and extended redshift coverage of DESI DR2 sustained the preference for a dynamical dark energy component, yielding a level of statistical significance comparable to, or greater than, that reported with DESI DR1 DESI:2025wyn ; DESI:2025fii ; DESI:2025zgx . These hints may have enormous implications for the understanding of the nature of DE, and align with expectations from certain fundamental theory frameworks Brandenberger:2025hof .
@@
In light of these results, recent works, among them Jiang:2024xnu and Escamilla:2024fzq , have pointed to signals, from late-time cosmological data, of oscillatory/non-monotonic features in the shape of the expansion rate at low redshifts  [^6] See, however, Akarsu:2022lhx for a warning that oscillatory features in $H(z)$ can be induced when fitting for this function given incomplete data. Also, the DESI full shape modelling does not show a signal Colgain:2024mtg .. The possibility of an oscillating dark energy has already been addressed in earlier works, as for instance in Refs. Zhao:2017cud ; Tamayo:2019gqj ; Rubano:2003er ; Linder:2005dw ; Feng:2004ff ; Nojiri:2006ww ; Kurek:2007bu ; Jain:2007fa ; Saez-Gomez:2008mkj ; Kurek:2008qt ; Pace:2011kb ; Pan:2017zoh ; Panotopoulos:2018sso ; Rezaei:2019roe ; Yao:2022jrw ; Rezaei:2024vtg ; Tian:2019enx ; DESI:2024kob ; Adil:2023exv ; Mbonye:2022cnf ; Mbonye:2024mss . In Ref. Escamilla:2024fzq , in light of the recent cosmological data, including DR1, twelve different parametrizations for an oscillatory dark energy equation of state were analyzed, among them the ones proposed in Refs. Linder:2005dw ; Feng:2004ff ; Pan:2017zoh ; Zhao:2005vj . This analysis demonstrated that all the oscillating DE parametrizations considered provide a better fit to the DR1 when compared to the cosmological constant, although in the Bayesian evidence analysis they are penalized due to the additional free parameter. In a similar context, in the work of Ref. Jiang:2024xnu , the late-time expansion history was reconstructed in a way, according to the authors, to be “as model-independent and non-parametric as possible” (see also Heisenberg:2020ywd ; Haude:2019qms ). In their first reconstruction, using DR1 + PantheonPlus data, two features in the unnormalized expansion rate were identified, a bump at redshift $z\sim 0.5$, and a depression at $z\sim 0.9$ (both features were absent when replacing the DESI dataset with the older SDSS one). On the other hand, the depression feature was noticeable in both the DR1+PantheonPlus and DR1+Union3 reconstructions, but not in the DR1+DESY5 one  [^7] For more information on the reconstruction methods see Ref.Jiang:2024xnu .. While the Chevallier-Polarski-Linder (CPL) fit, the $w_{0}-w_{a}$ parametrization often considered when interpreting DESI data, is able to partially capture the data characteristic behaviour, it cannot fit all the features, since the model does not have sufficient complexity. Achieving this may require an oscillatory/non-monotonic behavior of the Hubble expansion rate Jiang:2024xnu . More recently, some works (see Kessler:2025kju for instance) found evidence for an oscillating behaviour of the dark energy equation of state also from the more recent DR2 data.
@@
While a more conclusive result will need to wait for more precise future data, we believe it is now important to further investigate theoretical mechanisms that could account for a non-monotonic dynamical behavior of DE. As discussed above, the back-reaction effect of super-Hubble fluctuations is one mechanism that could naturally lead to an oscillatory behavior for the dark energy density, and consequently for the Hubble parameter, without the need for extra ingredients beyond the standard $\Lambda$CDM model. This back-reaction effect is present in any cosmological model whenever super-Hubble cosmological fluctuations are present.
@@
In the present work we propose a parametrization to account for the back-reaction effect of super-Hubble fluctuations by means of an oscillating effective energy density component. This may provide a full non-perturbative description of the back-reaction effect. We then analyze such a model considering data from the radial component of the Baryon Acoustic Oscillations from the recent DESI DR2 data. We compute the Hubble parameter at low redshifts as predicted by this model, and compare the result with the data in order to place constraints on the maximum allowed amplitude of the oscillations. We also analyze the predictions of the model for the effective dark energy equation of state, for the deceleration parameter, and for $H_{0}$.
@@
Our study can be considered to be a first-step analysis limited to background quantities, aiming only to highlight the importance of not neglecting the back-reaction mechanism. Nevertheless, in order to obtain more conclusive results, it is necessary to further analyze the predictions of the model, in particular taking into account the induced cosmological perturbations. This will allow us to test the model using additional BAO data beyond only the radial component, as well as using further data from other cosmological sources. This analysis is left for a future work.
@@ II A Model for the Back-Reaction of Super-Hubble fluctuations
In the absence of anisotropic stress, the scalar metric fluctuation variables $\phi$ and $\psi$ in eq.2 are equal. To write the metric, we have chosen a particular coordinate system (longitudinal gauge)  [^8] See e.g. Mukhanov:1990me for a review of the theory of cosmological perturbations.. In Eq. 2 $\gamma_{ij}$ is the background metric of the constant time hypersurfaces. We assume vanishing spatial curvature and then $\gamma_{ij}=\delta_{ij}$. In the above equations, the background metric is given by the scale factor $a(t)$ and the background matter by $\varphi_{0}(t)$. We will only consider here the effects of scalar metric fluctuations.
@@
where the first term on the right hand side of the equation is the contribution of the background matter field to $T_{\mu\nu}$. Improved studies of the back-reaction of super-Hubble fluctuations focus on physically measurable quantities such as the spatial average of the local Hubble expansion rate Geshnizjani:2003cn ; Marozzi:2012tp ; Finelli:2011cw ; Gasperini:2009wp ; Gasperini:2009mu ; Marozzi:2010qz ; Marozzi:2011zb ; Abramo:2001dd ; Losic:2005vg ; Losic:2006ht ; Brandenberger:2004ix ; Afshordi:2000nr even beyond perturbation theory Brandenberger:2018fdd . Some of these results have pointed to a self adjusting mechanism for the cosmological constant, as first discussed in Brandenberger:1999su .
The origin of the self-adjustment mechanism for the relaxation of the cosmological constant is as follows Brandenberger:1999su .: in an accelerating phase, fluctuation modes exit the Hubble radius and freeze out. Cosmological perturbations induced by super-Hubble matter fluctuations contribute to the effective energy-momentum tensor which has the form of a negative cosmological constant (negligible kinetic and tension energies leading to an effective equation of state $w=-1$, and negative effective energy density - negative since matter induces a potential well and the negativity of the gravitational energy overwhelms the positivity of the matter energy for super-Hubble modes). This will induce a decrease in the effective cosmological constant by counteracting the effect of the pre-existing $\Lambda$. But once the energy density $\rho_{DE}(t)$ in the effective cosmological constant has fallen below that of matter, accelerated expansion will stop, modes will re-enter the Hubble radius, the decline in $\rho_{DE}(t)$ halts, and within a fraction of the Hubble time, a renewed buildup of the backreaction effect takes place.
@@
where $A$ is a dimensionless amplitude, $\omega$ is the frequency of the oscillation, $\phi$ is a phase and $\rho_{m}$ is the background density of matter. The term $\omega t$ in the above equation can be written in terms of the redshift $z$ as $\omega t\equiv nz$, with $n$ being a constant, which we expect to have an order of magnitude not much smaller or bigger than one, since the typical time scale of oscillations is expected to be set by the Hubble time. This ansatz expresses the idea that, after the time $t_{eq}$ of equal matter and radiation, the back-reaction term should track the energy density $\rho_{m}$ of matter 999 Note that an oscillatory contribution to the dark energy density was recently also shown to emerge from a model in which the dark energy field emerges from the dynamics of a highly curved field space motivated by string theory Payeur:2024kyy .
A priori it is expected that the amplitude $A(t)$ should be time-dependent. However, in the present work, we will focus on the late time predictions, since we are going to analyze our model in light of data in the range $2.50\leq z<$ 2.5. For this small redshift interval we believe it is reasonable to consider $A(t)$ to be constant. Therefore, in the rest of the paper we are going to consider the following expression for the effective density from the back-reaction effect
$\rho_{BR}\ =\,A\;cos(nz+\phi)\;\rho_{m}(t)\,.$ (8)
The contribution of the energy density from the back-reaction effect is added to the total energy of the universe, in such a way that the first Friedmann equation can be written as,
@@
Concerning the phase $\phi$, we will choose it in a way that the back-reaction mechanism begins at a redshift $z=3400$ with $\rho_{BR}=0$, when it starts building up. This choice is reasonable since we are taking the back-reaction energy density as being proportional to the matter energy density, and before recombination it would rather track the radiation component.
@@
$\begin{split}w_{BR}&=\left(\frac{H^{\prime}H}{a}-\frac{3H^{2}}{2}+\frac{3H_{0}^{2}\;\Omega_{\Lambda 0}}{2}\right)\frac{1}{4\pi G}\;\frac{1}{\rho_{BR}},\\ &=-\frac{n\left(z+1\right)}{3\tan\left(n(z-3400)\right)}\end{split}$ (14)